NoFly Corporation sells three different models of a mosquito “zapper.” Model A12 sells for $60 and has variable costs of $41. Model B22 sells for $108 and has variable costs of $76. Model C124 sells for $405 and has variable costs of $313. The sales mix of the three models is A12, 58%; B22, 30%; and C124, 12%. If the company has fixed costs of $234,284, how many units of each model must the company sell in order to break even? (Round Per unit values to 2 decimal palces, e.g. 15.25 and final answers to 0 decimal places, e.g. 5,275.) Model A12 B22 C124 Total break-even units
Answer #1
Model A12 Model B22 Model C124 Total
Selling price 60 108 405
Less: Variable costs 41 76 313
Unit Contribution margin 19 32 92
X Sales mix percentage 58% 30% 12%
Weighted Unit Contribution margin 11.02 9.60 11.04 31.66
Fixed costs 234284
/ Weighted Unit Contribution margin 31.66
Overall break even 7400
Break even units:
Model A12 4292 =7400*58%
Model B22 2220 =7400*30%
Model C124 888 =7400*12%
Total break-even 7400 units